# KKL Observer Synthesis for Nonlinear Systems via Physics-Informed Learning

### v1

@@ 9 Conclusion
We proposed a learning method to design KKL observers for autonomous nonlinear systems. We trained neural networks using physics-informed learning to approximate both the transformation map and its inverse, which are required for synthesizing a KKL observer. By fixing the KKL observer, we synthetically generated the training data for neural networks by numerically solving system and observer equations. Instead of simultaneously learning the transformation and its inverse, we proposed to learn these maps sequentially to avoid conflicting gradients during training. First, the transformation map is learned by a physics-informed neural network regularized by the PDE associated with the KKL observer. Then, using the learned transformation map, its inverse is learned by a neural network with an autoencoder-type architecture.
We provided theoretical robustness guarantees and showed that the state estimation error obtained by the learned KKL observer is input-to-state stable against approximation errors and system uncertainties, i.e., model uncertainties and measurement noise. In addition, we presented a method for optimal KKL observer parameter selection that achieves a trade-off between robustness and learnability. The effectiveness of our approach was validated through comprehensive state estimation experiments across multiple benchmark systems: the reverse Duffing oscillator, van der Pol oscillator, Rössler attractor, and Lorenz attractor. Furthermore, we successfully demonstrated the learned KKL observer’s capability in fault detection and isolation within a network of Kuramoto oscillators.
Our future research will address key open problems. The primary challenge lies in extending KKL observer learning to non-autonomous and controlled nonlinear systems, where the transformation maps and their inverses are time-varying. This temporal dependency results in time-dependent PDEs that probably require more complex neural network architectures. Another challenge lies in incorporating our KKL observer learning framework into the design of output feedback controllers while maintaining stability guarantees.
@@
This work is supported by the Swedish Research Council and the Knut and Alice Wallenberg Foundation, Sweden. It has also received funding from the European Union’s Horizon Research and Innovation Programme under Marie Skłodowska-Curie grant agreement No. 101062523.

### v2

@@ 9 Conclusion
We proposed a learning-based method for designing KKL observers for autonomous nonlinear systems. We trained neural networks using synthetically generated data to approximate the transformation map and its inverse, which are required for synthesizing a KKL observer. Instead of simultaneously learning the transformation and its inverse, we proposed learning them sequentially to avoid conflicting gradients during training and improve accuracy. First, the transformation map is learned by a physics-informed neural network that incorporates the PDE associated with the KKL observer via automatic differentiation. Then, by generating new data and using the learned transformation map to produce corresponding features, a second neural network learns the inverse map by minimizing the empirical reconstruction risk.
Theoretically, we provided non-asymptotic generalization bounds for the learned KKL observer. We proved that minimizing the empirical risk guarantees the input-to-state stability of the state estimation error with respect to approximation errors and system uncertainties. The effectiveness of our approach was validated through comprehensive state estimation experiments across multiple benchmark systems. We demonstrated the generalization capability of the learned KKL observer by comparing it with the state-of-the-art methods. The excellent performance is due to that the PDE constraint serves as a physically consistent inductive bias that effectively reduces the hypothesis space for neural networks, enabling extrapolation where purely data-driven methods fail.
Our future research will address some key open problems. The primary challenge lies in extending KKL observer learning to non-autonomous and controlled nonlinear systems, where the transformation maps and their inverses are time-varying. This temporal dependency results in time-dependent PDEs that probably require more complex neural network architectures. Moreover, combining PINNs for system identification with observer design would address the unknown system model scenario more robustly than current end-to-end learning methods. Finally, we will also explore active learning to identify where in the state space the PDE residual is high and sample adaptively from those regions, further reducing the data requirement.
@@
This work is supported by the European Union’s Horizon Research and Innovation Programme under Marie Skłodowska-Curie grant agreement No. 101062523. It has also received funding from the Swedish Research Council’s Distinguished Professor Grant and the Knut and Alice Wallenberg Foundation’s Wallenberg Scholar Grant.